OpenStudy (anonymous):

f(x) = 16x5 – 48x4 – 8x3 and g(x) = 8x2. f(x)/g(x) A. 2x^3 + 6x^2 + x B. 2x^3 – 6x^2 – x C. 2x^2 + 6x + 1 D. 2x^2 – 6x – 1

OpenStudy (lalaly):

\[\large{\frac{16x^5 -48x^4-8x^3}{8x^2} = \frac{16x^5}{8x^2}-\frac{48x^4}{8x^2}-\frac{8x^3}{8x^2}}\]

OpenStudy (lalaly):

\[\large{2x^3-6x^2-x}\]